The distance between the first and the third nodes of a standing wave is 0.2 m, its maximum displacement is 0.02 m and its frequency is 40 Hz. The wave-functions of the two travelling waves which interfere to give the standing wave are then:


y1 = 0.02 sin⁡(20πx-320πt) ; y2 = 0.02 sin⁡(20πx+320πt),


y1 = 0.01 sin⁡(5πx-40πt) ; y2 = 0.01 sin⁡(5πx+40πt),


y1 = 0.01 sin⁡(10πx-80πt) ; y2 = 0.01 sin⁡(10πx+80πt),


y1 = 0.005 sin⁡(5πx-40πt) ; y2 = 0.005 sin⁡(5πx+40πt),


y1 = 0.04 sin⁡(20πx-320πt) ; y2 = 0.04 sin⁡(20πx+320πt),


1
Expert's answer
2021-05-24T09:02:35-0400

from the given information we know that:


maximum displacement =amplitude of standing wave=2A=0.02


"\omega" (angular frequency)="2\pi f" ="80\pi"

k="10\pi"


we get standing wave as:

"0.04sin(10\pi x)sin(80\pi t)"


using formula

2sinC SinD= Sin(C-D)+sin(C+D)


we get equations

(option C)

y1 = 0.01 sin⁡(10πx-80πt) ;

y2 = 0.01 sin⁡(10πx+80πt),